OpenStudy (anonymous):

Let f(x) = x2 - 8x + 5. Find f(-1)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

No :(

OpenStudy (skullpatrol):

$$\Huge f(x) = x^2 - 8x + 5$$$$\large f(-1)=(-1)^2-8(-1)+5$$

OpenStudy (sweetburger):

when you replace the x value in F(x) for another value you also replace each corresponding x value on the other side of the equation with that value example F(x)=2x^2+2x then you say F(-2) = 2(-2)^2+2(-2) and solve from there